Q: The annual bank discount rate on T-bill matures in 45 days, priced at $99.00 with a $100 face value, is ____:

A: Bank Disc Rate = 1.0% * 8 = 8.0%

Above is the question and the solution but, I am confused about how they came to this solution. I'm not sure where they got the 8 from? Any help would be appreciated!
Answer #1

Discount rate for 45 days =(100-99)/100=1%
The bank's discount rate =1%*360/Number of days =1%*360/45 =1%* 8 =8%